OpenStudy (anonymous):

F(x)=3x+1 and f-1=x-1/3, then f(2)=

OpenStudy (anonymous):

f(2)=3*2+1=6+1=7

OpenStudy (anonymous):

@keepsecure93

OpenStudy (anonymous):

Ok what happen to the negative.

OpenStudy (anonymous):

arcf=2-1/3=1/3

OpenStudy (anonymous):

f^-1=arcf

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

So in this one. F(x)=3x+1and f-x-1/3, then f-1(7). F(7)=3*7+1=21+1=22 is that right.

OpenStudy (anonymous):

Evaluate 3[5+3[10+4*8]]